[Brahms-l] Shift report 20050319 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Mar 19 2005 - 08:05:59 EST
Shifter leader was SB
Shifters were 


1. Two stores. The first one lasted 5 hours. The second one is still running.
2. Run No.  -- Fs@8 deg @A1723 -- MRS@45 deg @B1050
-------------------------------------------------
   13861        trig2(19k)        trig3(596k)
   13862        still running
-------------------------------------------------
3. No detector problems detected

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Mar 19 08:06:10 2005

This archive was generated by hypermail 2.1.8 : Sat Mar 19 2005 - 08:06:18 EST